### v2

@@ 4 Checklist-based Performance Assessment using Clear
In this section, we introduce Clear for expert-aligned evaluation of model performance on expert-level tasks. As shown in Fig. 1, given a model output, our framework maps its information to items in the checklist derived from the expert-design rubric, and assesses the quality of the model output by comparing the checklist-mapped model output against the checklist-mapped human reference (§4.1). We quantitatively justify key design choices in our evaluation pipeline, including the selection of the checklist-mapper and the judge for checklist comparison (§4.2).
@@ 4.1 Evaluation Process
We follow the same procedure described in §3.2 to extract checklists from model responses. However, instead of using GPT-4o, which can be significantly more expensive while extracting checklists from all models, we opt for the open-source model Qwen2.5-72B as the checklist mapper considering its availability of model weights and decent extraction performance, which will be validated in §4.2.
@@
To evaluate checklist-mapped responses, we assess the degree to which the checklist-mapped model response aligns with that of the reference. To this end, we use GPT-4o within an LLM-as-a-judge paradigm [24, 25, 26], adapting the reference-based scoring methodology [27] to evaluate the semantic alignment between each checklist item in a model response and the corresponding information in the reference for every sample. For instance, consider a checklist item $c$ with corresponding information in the model response and reference denoted as $H(c)$ and $R(c)$, respectively. The LLM judge assigns a binary score to each checklist item by evaluating whether the semantic content of $R(c)$ is contained within $H(c)$. The prompt for the judge is included in Appendix  LABEL:ap:eval-procedure. Despite suggestions to employ multiple models from different families as judges to mitigate biases [27], we use only GPT-4o, as it saves cost and shows high agreement with candidate judge models, as discussed in §4.2.
After performing item-level assessment for a checklist-mapped model response, we define its checklist precision (checklist recall) as the fraction of checklist items whose model response (reference) is semantically contained within the reference (model response) and accuracy as the fraction of checklist items whose model response and reference mutually contain each other. The sample-level checklist F1 score is the harmonic mean of the checklist precision and checklist recall, and we obtain the task-level performance by averaging the sample-level metrics. [^2] For simplicity, in the rest of the paper, we use precision, recall, and F1 score to indicate these checklist-specific metrics.
@@ 4.2 Evaluation Component Validation
Model Selection for Checklist Mapper. To enable accessible and low-cost checklist mapping, we evaluate open-weight models—Llama-3.3-70B-Instruct, Mistral-Large-Instruct, and Qwen2.5-72B —using reference checklists extracted by GPT-4o. We, first, validate the quality of these reference checklists through human and automated evaluations on tasks T1 and T6, confirming over 90% faithfulness and coverage. We selected these two tasks due to their challenging long contexts and extended output requirements involving over $25$ checklist items. To identify a suitable open-weight mapper, we evaluated tasks T1, T6, T7, and T8 which spans diverse domains and configurations of input / output length. Qwen2.5-72B achieved the highest average performance with an average F1-score of 90.1%, demonstrating its applicability for accurate checklist mapping. See Appendix LABEL:app:judge_analysis for details on above analyses.
@@ 5 Experiments
We evaluate 11 models from 3 open-weight families (Llama [28] 8B, 70B; Mistral 12B, 123B; and Qwen [29] 7B, 72B) and 3 proprietary families (GPT-4o, GPT-4o-mini; Claude-3.5-Haiku, Claude-3.7-Sonnet; and Gemini-2.0-Flash  [^3] We include only Gemini-2.0-Flash, as the Gemini-2.0 series lacks a variant comparable in size to GPT-4o or Claude-3.7-Sonnet. Gemini-2.5 models are excluded due to frequent empty outputs without explanation.) on ExpertLongBench. For open-weight models, we use their instruction fine-tuned and RLHF variants. Model details are provided in Appendix LABEL:app:experiment_details.
@@
Tab. 2 presents the performance of the evaluated models, measured with the F1 score. The models are sorted by their average F1 score across all tasks. Accuracy, precision, and recall are reported in LABEL:tab:main_accuracy_results, LABEL:tab:main_precision_results, and LABEL:tab:main_recall_results. The tasks in ExpertLongBench pose significant challenges to existing LLMs. Notably, Gemini-2.0-Flash, the top-performing model in our evaluation, achieves an average F1 score of only 26.8, underscoring the difficulty of the benchmark. Among all tasks, T2 proves to be the most challenging, with all models scoring an F1 below 8. In addition to the complex nature of legal argumentation, T2 involves longer inputs and necessitates longer outputs, thereby further testing the long-context capabilities of the models.
@@
We observe that within the same model family, the larger model outperforms their smaller counterpart in terms of average performance. However, this superiority is not consistent across all individual tasks. For example, on task T4, Mistral-Large-Instruct ranks 2nd while Mistral-Nemo-Instruct ranks 9th. Conversely, on task T10, Mistral-Large-Instruct ranks 6th, whereas Mistral-Nemo-Instruct achieves the best performance across all models. In fact, no single model consistently outperforms its smaller variant from the same family across all tasks. This inconsistency may suggest that the pre-training curricula of existing LLMs might not uniformly emphasize or adequately represent the domain-specific tasks, especially when scaling for larger models, potentially leading to an imbalance in their capabilities across different tasks.
@@
Figure 2: F1 score vs. coverage of checklist items. The coverage of checklist items is the percentage of checklist items that are covered in the generated text regardless their correctness.
@@
High coverage of checklist items does not imply high output quality.
We further investigate the percentage of checklist items covered in the generated text, irrespective of correctness, and its correlation with output quality, as measured by the F1 score. Fig. 2 illustrates the relationship between checklist item coverage and the F1 score, aggregated across all experimented models and tasks, as well as presented separately for task T10. A positive correlation between coverage and F1 score is observed for most tasks, resulting in an overall positive correlation when averaged. However, the correlation coefficient remains low—models frequently obtaining high coverage scores while concurrently exhibiting low F1 scores, suggesting their ability to produce content that seemingly adheres to domain-specific standards but incorporates incorrect information. Such behavior presents a risk, as the output might mislead users into perceiving the generated content as reliable due to its apparent completeness, despite its underlying inaccuracies. The issue is notably more pronounced for T10 (ESG report summarization), where coverage exhibits a negative correlation with output quality. The structure of ESG reports provides hints for elements that should be covered in the output. However, due to the volume of information, models cannot precisely identify which information is most critical to stakeholders. Further main results and detailed outcomes for each task can be found in Appendix LABEL:ap:additional_results.
@@ 6 Towards Reproducible and Low-Cost Evaluation
While GPT-4o demonstrates strong performance in evaluating checklist-aligned model responses [30, 31, 32, 33, 34], there are three primary concerns associated with the use of closed models: (a) they are typically accessible only through paid APIs, with costs that can escalate rapidly in large-scale experiments involving multiple queries or extended outputs; (b) their deployment on external servers raises data privacy and security concerns, particularly in sensitive domains such as healthcare, law, and education; and (c) proprietary models are subject to unannounced updates or version changes, which can compromise the reproducibility of results across studies or even within the same research group over time.
Therefore, the primary objective of this section is to investigate whether the judge component in our evaluation pipeline can be substituted with open-weight alternatives to promote a more accessible and reproducible research environment. To this end, we evaluate various open-weight models and their combinations as judge models by measuring how well their accuracy scores correlate with those from GPT-4o, which serves as the ground truth due to its strong domain-level performance. For model combinations involving multiple judge models, we explore two pooling strategies [35, 27] to compute an aggregate score for each sample: (a) Mean Pooling, where the final score is the average of the scores assigned by each constituent judge; and (b) Majority Pooling, where the final score corresponds to the mode of the individual scores. In cases where multiple modes exist, we resolve the tie by averaging their values.
@@
The results presented in LABEL:tab:main_correlation_results show the correlation among the computed scores of various model variants. In this table, the category labeled Small combines scores from Llama3.1-8B-Instruct, Mistral-Nemo-Instruct, and Mistral-Large-Instruct. Similarly, the category Large includes Llama-3.3-70B-Instruct, Mistral-Large-Instruct, and Qwen2.5-72B. The category All-small encompasses models from Small and GPT-4o-mini.
Best performing single model judge.
From Fig. 3, we find that most judge models exhibit a high degree of correlation with GPT-4o with Qwen2.5-72B yielding the highest average correlation among all single judges. This indicates that a single model can reliably assess checklists while significantly maintaining computational overhead to a low value.
@@
Additionally, combining open-source models enhances alignment with GPT-4o. For instance, while Llama3.1-8B-Instruct, Mistral-Nemo-Instruct, and Qwen2.5-7B individually achieve average correlations of 0.65, 0.80, and 0.78, respectively, majority pooling their outputs raises the correlation to 0.82. Moreover, in most model combinations—excluding large-open-source-models—majority pooling consistently outperforms mean pooling in terms of correlation. Since majority pooling is more effective for estimating binary variables [27], and our checklist scoring relies on binary labels, it leads to more accurate overall score assignment.
Tasks T1 and T2 exhibit the lowest average correlations as shown in LABEL:tab:main_correlation_results and weakest model performance referring to Tab. 2, indicating that higher task complexity may hinder reliable evaluation.
@@
A regression analysis of the relationship between task complexity and average judgment correlation with GPT-4o (see Fig. 4) reveals a strong correlation and a moderately high $R^{2}$ value, indicating that the regression model explains a substantial portion of the variance. This can be used to decide the scale of model for evaluating a particular task. For instance, smaller models like Qwen2.5-7B and Mistral-Nemo-Instruct align well with GPT-4o for less complex tasks. For more challenging tasks such as T1 and T2, larger models like Qwen2.5-72B are recommended.
@@ 7 Skill Decomposition Analysis
We provide a detailed skill-level and difficulty-level analysis of model performance on ExpertLongBench for a better understanding of where current models excel and where they fall short. Our analysis is grounded in a fine-grained examination of each task’s checklist items, identifying the specific skills required to fulfill each item at corresponding difficulty levels. Detailed explanations of the skills, difficulty levels, and item-level skill and difficulty mapping are provided in Appendix LABEL:app:skill_analysis.
We examine model performance across varying knowledge levels including Below College, College, and Graduate, and reasoning difficulty including Low (Knowledge Memorization), Medium (Knowledge Understanding), High (Knowledge Applying), Very High (Knowledge Creating) [36, 37]. We present our key findings as follows.
(a) Model performance across various levels of knowledge difficulty.
(b) Model performance across various levels of reasoning complexity.
Figure 5: Model performance across various difficulty levels.
@@
The consistent performance drop at the Graduate-level knowledge as shown in Fig. 5(a) highlights a substantial challenge for current models in mastering expert-level knowledge. The clear performance decline with higher reasoning difficulty as shown in Fig. 5(b) emphasizes the importance of assessing models beyond knowledge memorization and understanding, focusing on their ability to perform complex reasoning. Most models perform worse on low reasoning difficulty level (knowledge memorization) than medium reasoning difficulty level (knowledge understanding) because the expert-level knowledge in ExpertLongBench may be long-tail knowledge that rarely appears during training, limiting memorization performance, while the model can still leverage its reasoning and generalization abilities to approximate understanding [38]. Notably, ExpertLongBench is specifically designed to assess more advanced knowledge processing skills.
@@
Models show divergent strengths. For instance, Claude-3.7-Sonnet ranks first in the low reasoning level (knowledge memorization) but ranks relatively low in the very high reasoning level (knowledge creation) (Fig. 5(b)). These results emphasize the value of benchmark with diverse difficulty levels, which captures performance differences across varying knowledge and reasoning levels, revealing model strengths and weaknesses that might be overlooked in general benchmarks. Notably, ExpertLongBench is meticulously designed to provide a comprehensive evaluation across diverse difficulty levels, and we recommend the community to leverage ExpertLongBench for future model evaluation. Analysis results on more general skills and additional details are in Appendix LABEL:app:skill_analysis.
@@ 8 Conclusion
We introduce ExpertLongBench, a multi-domain expert-level benchmark. ExpertLongBench comprises 1050 samples, spanning 11 tasks across 9 distinct domains. These tasks, originating from realistic expert applications rather than typical question-answering scenarios, demand long-form outputs that adhere to domain-specific standards. Each task is accompanied by an expert-designed and validated rubric, which serve as evaluation guidelines. To evaluate long-form model outputs on these expert tasks, we design Clear, an evaluation framework that maps model outputs and references into checklists derived from the rubric and subsequently performs an item-by-item comparison of these checklists. This process yields a fine-grained, grounded evaluation that aligns with domain-specific requirements. We conduct experiments with 11 LLMs, revealing that ExpertLongBench pose a significant challenge for current LLMs. While these models can generate content that superficially matches required aspects, they frequently lack accuracy. Further analysis of Clear demonstrates the feasibility of substituting proprietary models with open-weight alternatives for the roles of checklist mapper and evaluator, facilitating low-cost, scalable benchmarking.
@@ Appendix A Benchmark Description
  • Checklist-mapped reference: A checklist-based, fine-grained evaluation rubric is designed for each task through collaboration with domain experts and by referencing established resources. A corresponding checklist-mapped reference is constructed for every sample, as detailed in §3.2 in the main paper.

### v3

@@ 4 Checklist-based Performance Assessment using Clear
In this section, we introduce Clear for expert-aligned evaluation of model performance. As shown in Fig.˜1, given a model output, our framework maps its information to items in the checklist derived from the expert-design rubric, and assesses the quality of the model output by comparing the checklist-mapped model output against the checklist-mapped human reference (§4.1). We quantitatively justify key design choices in Clear, including the selection of the checklist-mapper and the judge for checklist comparison (§4.2).
@@ 4.1 Evaluation Process
We follow the same procedure described in §3 to extract checklists from model responses. However, instead of using GPT-4o, which can be significantly more expensive while extracting checklists from all models, we opt for the open-weight model Qwen2.5-72B as the checklist mapper considering its availability of model weights and decent extraction performance, which will be validated in §4.2.
@@
To evaluate checklist-mapped responses, we assess the degree to which the checklist-mapped model response aligns with that of the reference. To this end, we use GPT-4o within an LLM-as-a-judge paradigm (Gu et al., 2024; Li et al., 2024a, b), adapting the reference-based scoring methodology (Verga et al., 2024) to evaluate the semantic alignment between each checklist item in a model response and the corresponding information in the reference for every sample. For instance, consider a checklist item $c$ with corresponding information in the model response and reference denoted as $H(c)$ and $R(c)$, respectively. The LLM judge assigns a binary score to each checklist item by evaluating whether the semantic content of $R(c)$ is contained within $H(c)$. In other words, we assign a binary score of 1 only when all the information conveyed by $R(c)$ is also present in $H(c)$. The prompt for the judge is in Appendix C. We use GPT-4o as the judge, as it saves cost and shows high agreement with candidate judge models, as discussed in §4.2.
After performing item-level assessment for a checklist-mapped model response, we define its checklist precision (checklist recall) as the fraction of checklist items whose model response (reference) is semantically contained within the reference (model response) and accuracy as the fraction of checklist items whose model response and reference mutually contain each other. We obtain the task-level performance by averaging the sample-level metrics.
@@ 4.2 Evaluation Component Validation
Model Selection for Checklist Mapper. To enable cost-effective and reproducible checklist mapping, we primarily consider open-weight models—Llama-3.3-70B-Instruct, Mistral-Large-Instruct, and Qwen2.5-72B —and evaluate them using the reference checklists extracted by GPT-4o. We, first, validate the quality of these reference checklists through human and automated evaluations on tasks T1 and T6, confirming over 90% faithfulness and coverage. We selected these two tasks due to their challenging long contexts and extended output requirements involving over $25$ checklist items. To identify a suitable open-weight mapper, we evaluated tasks T1, T6, T7, and T8 which spans diverse domains and configurations of input / output length. Qwen2.5-72B achieved the highest average performance with an average F1-score of 90.1, demonstrating its applicability for accurate checklist mapping. [^5] To investigate more cost-efficient options for checklist mapping, we also evaluate smaller models from the same family as Qwen2.5-72B, as detailed in Appendix E. While these models achieve reasonable performance, they still fall notably short of Qwen2.5-72B. Consequently, we continue to use Qwen2.5-72B as the checklist mapper to ensure high-quality and accurate mappings. See Appendix E for details on above analyses.
@@ 5 Experiments
We evaluate 13 models from 3 open-weight families and 3 proprietary families on ExpertLongBench. For open-weight models, we use their instruction fine-tuned and RLHF variants. Model details are provided in Appendix G.
@@
Tab.˜2 presents the performance of the evaluated models, measured with the F1 score. The models are sorted by their average F1 score across all tasks. Accuracy, precision, and recall are reported in Tab.˜45, Tab.˜46, and Tab.˜47. The tasks in ExpertLongBench pose significant challenges to existing LLMs. Notably, Gemini-2.5-Pro, the top-performing model in our evaluation, achieves an average F1 score of only 33.4, underscoring the difficulty of the benchmark. This also indicates substantial room for improvement, highlighting the benchmark’s longevity and its continued value as a challenging resource for evaluating future LLMs. Among all tasks, T2 proves to be the most challenging, with all models scoring an F1 below 11. In addition to the complex nature of legal argumentation, T2 involves longer inputs and necessitates longer outputs, thereby further testing the long-context capabilities of the models.
@@
We observe that within the same model family, the larger model outperforms their smaller counterpart in terms of average performance. However, this superiority is not consistent across all individual tasks. For example, on T4, Mistral-Large-Instruct ranks 2nd while Mistral-Nemo-Instruct ranks 11th. Conversely, on task T10, Mistral-Large-Instruct ranks 6th, whereas Mistral-Nemo-Instruct achieves the best performance across all models.
Figure 2: F1 score vs. coverage of checklist items (i.e., the percentage of checklist items that are covered in the generation regardless their correctness).
In fact, no single model consistently outperforms its smaller variant from the same family across all tasks. This inconsistency may suggest that the pre-training curricula of existing LLMs might not uniformly emphasize or adequately represent the domain-specific tasks, especially when scaling for larger models, potentially leading to an imbalance in their capabilities across different tasks.
@@
High coverage of checklist items does not correlate with quality.
We further investigate the percentage of checklist items covered in the generated text, irrespective of correctness, and its correlation with output quality, as measured by the F1 score. Fig.˜2 illustrates a overall negative correlation between checklist item coverage and the F1 score, aggregated across all experimented models and tasks. Models frequently obtaining high coverage scores while concurrently exhibiting low F1 scores, suggesting their ability to produce content that seemingly adheres to domain-specific standards but is incorrect. Such behavior presents a risk, as the output might mislead users into perceiving the generated content as reliable due to its apparent completeness Further main results and detailed outcomes for each task can be found in Appendix D.
@@ 6 Towards Reproducible and Low-Cost Evaluation
While GPT-4o demonstrates strong performance in evaluating checklist-aligned model responses (Xu et al., 2024; Posner and Saran, 2025; Seßler et al., 2025; Jo et al., 2024; Eriksen et al., 2024), there are three primary concerns associated with the use of closed models: (a) Performing large-scale experiments are costly, (b) Their deployment on external servers poses privacy risks in sensitive domains, and (c) Unannounced updates or version changes that undermine reproducibility across studies.
@@
Therefore, the primary objective of this section is to investigate whether the judge component in our evaluation pipeline can be substituted with open-weight alternatives to promote a more accessible and reproducible research environment. To this end, we evaluate various open-weight models and their combinations as judge models by measuring how well their accuracy scores correlate with those from GPT-4o, which serves as the ground truth due to its strong domain-level performance. For model combinations involving multiple judge models, we explore two pooling strategies (Wang et al., 2022; Verga et al., 2024) to compute an aggregate score for each sample: (a) Mean Pooling, where the final score is the average of the scores assigned by each constituent judge; and (b) Majority Pooling, where the final score corresponds to the mode of the individual scores. In cases where multiple modes exist, we resolve the tie by averaging their values.
The results presented in Tab.˜51 show the correlation among the computed scores of various model variants. In this table, the category labeled Small combines scores from Llama3.1-8B-Instruct, Mistral-Nemo-Instruct, and Mistral-Large-Instruct. Similarly, the category Large includes Llama-3.3-70B-Instruct, Mistral-Large-Instruct, and Qwen2.5-72B. The category All-small encompasses models from Small and GPT-4o-mini.
Performance of single-model judges.
From Fig.˜3, we find that most judge models exhibit a high degree of correlation with GPT-4o, with Qwen2.5-72B yielding the highest average correlation among all single judges. This indicates that a single model can reliably assess checklists while significantly maintaining computational overhead to a low value.
@@
Additionally, combining open-source models enhances alignment with GPT-4o. For instance, while Llama3.1-8B-Instruct, Mistral-Nemo-Instruct, and Qwen2.5-7B individually achieve average correlations of 0.65, 0.80, and 0.78, respectively, majority pooling their outputs raises the correlation to 0.82. Moreover, in most model combinations—excluding large-open-source-models—majority pooling consistently outperforms mean pooling in terms of correlation.
@@
A regression analysis of the relationship between task complexity and average judgment correlation with GPT-4o (see Fig.˜4) reveals a strong correlation and a moderately high $R^{2}$ value, indicating that the regression model explains a substantial portion of the variance. This can be used to decide the scale of model for evaluating a particular task. For instance, smaller models like Qwen2.5-7B and Mistral-Nemo-Instruct align well with GPT-4o for less complex tasks. For more challenging tasks such as T1 and T2, larger models like Qwen2.5-72B are recommended. Since majority pooling is more effective for estimating binary variables (Verga et al., 2024), and our checklist scoring relies on binary labels, it leads to more accurate overall score assignment.
@@ 7 Skill Decomposition Analysis
We provide a detailed skill-level and difficulty-level analysis of model performance on ExpertLongBench for a better understanding of where current models excel and where they fall short. Our analysis is grounded in a fine-grained examination of each task’s checklist items, identifying the specific skills required to fulfill each item at corresponding difficulty levels. Detailed explanations of the skills, difficulty levels, and item-level skill and difficulty mapping are provided in Appendix F.
We examine model performance across varying knowledge levels including Below College, College, and Graduate, and reasoning difficulty including Low (Knowledge Memorization), Medium (Knowledge Understanding), High (Knowledge Applying), Very High (Knowledge Creating) (Krathwohl, 2002; Yu et al., 2024). We present our key findings as follows.
Figure 5: Model performance across various levels of reasoning complexity.
@@
The consistent performance drop at the Graduate-level knowledge as shown in Fig.˜11 highlights a substantial challenge for current models in mastering expert-level knowledge. The clear performance decline with higher reasoning difficulty as shown in Fig.˜5 emphasizes the importance of assessing models beyond knowledge memorization and understanding, focusing on their ability to perform complex reasoning. Many models perform worse on low reasoning difficulty level (knowledge memorization) than medium reasoning difficulty level (knowledge understanding) because the expert-level knowledge in ExpertLongBench may be long-tail knowledge that rarely appears during training, limiting memorization performance, while the model can still leverage its reasoning and generalization abilities to approximate understanding (Kandpal et al., 2023). Notably, ExpertLongBench is specifically designed to assess more advanced knowledge processing skills.
@@
Models show divergent strengths. For instance, Claude-3.7-Sonnet ranks first in the low reasoning level (knowledge memorization) but ranks relatively low in the very high reasoning level (knowledge creation) (Fig.˜5). These results emphasize the value of benchmark with diverse difficulty levels, which captures performance differences across varying knowledge and reasoning levels, revealing model strengths and weaknesses that might be overlooked in general benchmarks. Notably, ExpertLongBench is meticulously designed to provide a comprehensive evaluation across diverse difficulty levels, and we recommend the community to leverage ExpertLongBench for future model evaluation. Analysis results on more general skills and additional details are in Appendix F.
@@ 8 Conclusion
We introduce ExpertLongBench, a multi-domain expert-level benchmark. ExpertLongBench comprises 1050 samples, spanning 11 tasks across 9 distinct domains. These tasks, originating from realistic expert applications rather than typical question-answering scenarios, demand long-form outputs that adhere to domain-specific standards. Each task is accompanied by an expert-designed and validated rubric, which serve as evaluation guidelines. To evaluate long-form model outputs on these expert tasks, we design Clear, an evaluation framework that maps model outputs and references into checklists derived from the rubric and subsequently performs an item-by-item comparison of these checklists. This process yields a fine-grained, grounded evaluation that aligns with domain-specific requirements. We conduct experiments with 13 LLMs, revealing that ExpertLongBench pose a significant challenge for current LLMs. While these models can generate content that superficially matches required aspects, they frequently lack accuracy. Further analysis of Clear demonstrates the feasibility of substituting proprietary models with open-weight alternatives for the roles of checklist mapper and evaluator, facilitating low-cost, reproducible and scalable benchmarking.
@@ Appendix A Benchmark Description
  • Checklist-mapped reference: A checklist-based, fine-grained evaluation rubric is designed for each task through collaboration with domain experts and by referencing established resources. A corresponding checklist-mapped reference is constructed for every sample, as detailed in §3 in the main paper.
@@
Experts Involved in Rubric Design and Verification.
We detail the experts involved in rubric design and verification as follows:
  • T1LegalMDS: We collaborated with a JD student at an R1 university in the US who created the checklist based on an established instruction for writing high-quality legal summaries. The JD student was compensated through supervisor funding support.
  • T2LegalSFG: This task involved a different set of experts from T1—two JD students and one legal professor from an R1 university in the US. The JD students were also compensated via supervisor funding.
  • T3MaterialSEG: The checklist was developed with input from a senior PhD student at an R1 university in the US specializing in material science, with extensive academic training and multiple publications, supporting the representativeness of the checklist.
  • T4EduPAE: We adopted the checklist from an existing, validated publication, which was confirmed by a professor in learning sciences at an R1 university in the US with a strong publication record.
  • T5EduFG: The checklist was adapted from instructional materials developed and taught over multiple years in the Econ 101 course at an R1 university in the US by a senior instructor.
  • T7ChemMDG: The checklist was developed in collaboration with a PhD student at an R1 university in the US who holds a master’s degree and has extensive academic training along with multiple publications.
  • T8BioPDG: Two graduate students studying biology at an R1 university in the US collaboratively created the rubric, and they reached consensus.
  • T10FinanceESG: We consulted professionals at a leading ESG rating institution. The checklist guidelines were taken and adapted from their ESG rating methodology document and the underlying key metrics.
  • T6HealthCNG, T9MedicalDR and T11CyberRDG: Checklists for these tasks were generated by the authors based on careful review of prior work and established relevant guidelines.